OpenStudy (anonymous):

solve 5/8x+1/16x=7/16. (I got x=-7/5)

OpenStudy (bobbyleary):

Start by finding a common denominator. In this case, 16 seems to be the easiest. \[(10/16)x + (1/16)x = (7/16)\] Now add common terms, so all the terms with x's. \[(11/16)x = (7/16)\] Now to solve for x, multiply (7/16) by (16/11). This results in: \[x = (7/11)\] Now to double check, plug x = 7/11 into the original problem, and it does work out.